# Closed-form finite-time blow-up and conditional perturbative stability framework for a (1+2)(1+2)D system (E1) derived from the 2D inviscid Boussinesq equations

### 31 Mar 2026 v2

@@ Abstract.
In polar variables on a planar wedge, we study a closed (1+2)(1+2)D system (E1) derived from the two-dimensional inviscid Boussinesq equations under a parity ansatz. The system retains a stretching-type interaction between scalar building blocks while remaining explicit enough to support both closed-form blow-up profiles and a weighted remainder analysis.
The paper has three main outputs. First, we identify the special ridge rays on which (E1) reduces to a convection-free (1+1)(1+1)D reaction system of Constantin–Lax–Majda type, and we record the resulting explicit finite-time blow-up mechanism. Second, after embedding these ridge profiles into the full wedge by anisotropic seeds, we derive the exact remainder equations around a prescribed background and prove singular weighted linear estimates for the remainder system. Third, we formulate a conditional nonlinear control principle in the spirit of Elgindi–Jeong: if the background exists on $[0,T)$ with the coefficient bounds required by the weighted energy method, and if the remainder remains subordinate to the background singularity in the detecting norm, then the full solution inherits the same finite-time blow-up.
What is unconditional in the present revision is the exact reduction from Boussinesq, the explicit ridge blow-up mechanism, and the weighted remainder framework. What remains conditional is the full-wedge background construction up to the blow-up time together with the rigidity properties needed to close the bootstrap without loss. In this sense, the manuscript isolates the background existence problem as the main remaining step toward a complete nonlinear stability theorem for the blow-up scenario.
@@ 1. Introduction
The question of whether smooth solutions can develop a singularity in finite time is a central theme in nonlinear PDE and fluid mechanics. In this paper we study a $(1+2)$-dimensional closed subsystem (E1), rigorously derived from the inviscid 2D Boussinesq equations: it retains a stretching-like interaction between two scalar fields while suppressing enough geometric and nonlocal complexity to permit closed-form analysis. Our aim is twofold: to construct explicit finite-time blow-up backgrounds and to formulate a perturbative stability theory around those backgrounds that is mathematically solid at the linear level and explicit about the remaining nonlinear obstruction.
@@
As emphasized by Elgindi–Jeong [17] and Elgindi–Pasqualotto [19], singularity formation for the inviscid 2D Boussinesq and 3D Euler systems has a long history. We recall only a small selection of representative references here, emphasizing works most closely aligned with the closed-form mechanism and stability framework developed below. For the inviscid 2D Boussinesq system, classical local theory and conditional breakdown criteria go back to Cannon–DiBenedetto [12], Chae–Nam [3] (see also Chae–Kim–Nam [7]), and Taniuchi [29]; see also Wu’s lecture notes [30] and the Euler perspective of Beale–Kato–Majda [1] and Constantin [10, 11]. On the modeling side, explicit/didactic mechanisms include Chae–Constantin–Wu [2] and the CLM/De Gregorio lineage [9, 16], as well as 1D reductions for Boussinesq-type dynamics such as Choi–Kiselev–Yao [8]. For rigorous singularity constructions and perturbative stability scenarios in PDE settings, see Elgindi–Jeong [18, 17], Chen–Hou [4, 5], and the synthesis of Drivas–Elgindi [15].
@@
    We identify the ridge rays on which (E1) reduces to a convection-free (1+1)(1+1)D reaction system of Constantin–Lax–Majda type and record the resulting explicit finite-time blow-up profiles.
@@
    We derive the exact remainder equations around a prescribed background in the updated $(x,\xi)$ variables, with all pure-background contributions retained in the background system.
@@
    We prove weighted singular linear estimates and formulate a conditional nonlinear remainder theorem of Elgindi type: once the background is available with the required coefficient bounds, blow-up transfers from the background to the full solution.
@@
    We isolate the remaining open step in the program, namely the full-wedge background construction up to the blow-up time together with the compatibility structure needed to close the nonlinear bootstrap.
Organization. Section 2 derives the closed subsystem (2.5) from the inviscid 2D Boussinesq equations and identifies the ridge rays on which (E1) reduces to a (1+1)(1+1)D CLM-type reaction system. Section 2.6 rewrites the background and remainder system in the updated $(x,\xi)$ variables and records the exact remainder equations. Section 4 collects the explicit background formulas and the corresponding weighted pointwise and energy bounds. Section 5 proves the singular weighted linear estimates, formulates the conditional nonlinear remainder mechanism, and explains how blow-up transfers from the background to the full solution once the remaining background hypotheses are verified.
@@ 2.4. Background
$\left\{\begin{aligned} \lambda U_{t}-\tfrac{1}{2}V\,U&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xU_{x}-(G+V)\xi U_{\xi},\\[5.69054pt] \lambda V_{t}-V^{2}+U^{2}&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xV_{x}-(G+V)\xi V_{\xi}+\tfrac{1}{x}P_{x}-\tfrac{1+\xi^{2}}{x^{2}}\xi{P_{\xi}},\\[5.69054pt] \lambda G_{t}+G^{2}&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xG_{x}-\tfrac{1}{2}\xi(G^{2}+V^{2})_{\xi}-\tfrac{\mu}{x}P_{x}-\tfrac{\mu(1+\xi^{2})}{x^{2}\xi^{2}}\xi{P_{\xi}},\\[5.69054pt] G-V&=-\tfrac{1}{1+\xi^{2}}(\xi^{2}xG_{x}-xV_{x})+\xi(G_{\xi}+V_{\xi}).\end{aligned}\right$. (2.9)
@@ 2.5. Ridge ray and Ridge functions
$\lambda=\tfrac{3}{2},\quad\mu=\left(\tfrac{5}{3}\right)^{1/2},$ (2.12)
@@
Separating $P_{x}$ from $V_{t}$ yields  (2.13) together with the simplified Poisson equation for $P(t,x,\xi_{0})$.
$P_{x}$ $=\tfrac{3x}{8}\left(U^{2}-2V^{2}\right),$ (2.16)
This proves Claim (C) and completes the Proof of theorem 2.5. ∎
@@
$\left\{\begin{aligned} &r=x^{2}+A_{1}(\xi^{2}-1)^{2m},\\ &a=a(x)=\frac{A}{\left(1+r\right)^{3}},\quad A>0,\\ &b=b(x)=\frac{Br}{\left(1+r\right)^{6}},\quad B>0.\end{aligned}\right$. (2.18)
@@ 2.6. Ridge blow-up and the full-wedge extension problem
The preceding theorem gives an explicit blowing-up ridge core. The full background problem is to extend that ridge core to a classical wedge solution of (2.9). The next proposition isolates the exact compatibility requirement behind this extension.
Theorem 2.7 (Compatibility condition for preserving ridge flatness).
@@
Then the first normal derivative of $U$ evolves by
$\lambda\,\partial_{t}(U_{\xi})(t,x,\xi_{0})=-\xi_{0}V(t,x,\xi_{0})\Bigl(xU_{x}(t,x,\xi_{0})+2U_{\xi\xi}(t,x,\xi_{0})\Bigr).$ (2.21)
Consequently, exact first-order ridge flatness is not automatically propagated by (2.9). In particular, for the seed family
@@
$\left\{\begin{aligned} &r=x^{2}+A_{1}(\xi^{2}-1)^{2m},\\ &a(x,\xi)=\frac{A}{\left(1+r\right)^{3}},\quad A>0,\\ &b(x,\xi)=\frac{Br}{\left(1+r\right)^{6}},\quad B>0,\end{aligned}\right$. (2.22)
with $m\geq 2$, one has $U_{\xi\xi}(0,x,\xi_{0})=0$; hence (2.21) reduces initially to
$\lambda\,\partial_{t}(U_{\xi})(0,x,\xi_{0})=-\xi_{0}V(0,x,\xi_{0})\,xU_{x}(0,x,\xi_{0}),$
which is generically nonzero for $x\neq 0$.
Proof.
Differentiate the $U$-equation in (2.9) with respect to $\xi$. Writing
$A(t,x,\xi):=\frac{\xi^{2}G-V}{1+\xi^{2}},\qquad B(t,x,\xi):=(G+V)\xi,$
the $U$-equation becomes
$\lambda U_{t}-\tfrac{1}{2}VU=-A\,xU_{x}-B\,U_{\xi}.$
Differentiating gives
$\lambda(U_{\xi})_{t}-\tfrac{1}{2}V_{\xi}U-\tfrac{1}{2}VU_{\xi}=-A_{\xi}xU_{x}-AxU_{x\xi}-B_{\xi}U_{\xi}-BU_{\xi\xi}.$
Now evaluate at $\xi=\xi_{0}$ under (2.20). Since $\xi_{0}^{2}=1$, $G=V$, and $G_{\xi}=V_{\xi}=0$ there, we have
$A(t,x,\xi_{0})=0,\qquad A_{\xi}(t,x,\xi_{0})=\xi_{0}V(t,x,\xi_{0}),\qquad B(t,x,\xi_{0})=2\xi_{0}V(t,x,\xi_{0}).$
Using also $U_{\xi}=V_{\xi}=0$ on the ridge, we obtain exactly (2.21). Finally, for the seed family (2.22) with $m\geq 2$, the dependence on $\xi$ enters through $(\xi^{2}-1)^{2m}$, so the initial second $\xi$-derivative vanishes at $\xi=\pm 1$. The last assertion follows. ∎
Remark 2.8 (What revision (c) actually proves).
Revision (c) establishes an explicit blowing-up ridge background and pinpoints the compatibility condition needed to extend it to a full wedge background. The ridge core itself is completely explicit by Theorem 2.6. What remains open is the full-wedge extension problem: one must construct a classical solution of (2.9) on $0,T)\times\Omega$ whose ridge trace agrees with that explicit core and whose normal derivatives satisfy the compatibility relation (2.21) (and the analogous relations for $V,G,P$).
In particular, the previous heuristic claim that a sufficiently large power $m$ in (2.22) should by itself preserve exact ridge flatness is too optimistic. Large $m$ still improves angular flatness of the initial seed and is useful for weighted estimates, but Proposition 2.7 shows that exact flatness is not an invariant of the full background flow without additional structure. This is now the main analytic issue on the background side of the manuscript.

### 03 Apr 2026 v3

# A pressure–velocity apex blow-up mechanism and conditional perturbative stability framework for a (1+2)(1+2)D system (E1) derived from the 2D inviscid Boussinesq equations

@@ Abstract.
In polar variables on a planar wedge, we study a closed (1+2)(1+2)D system (E1) derived from the two-dimensional inviscid Boussinesq equations under a parity ansatz. A central feature of the paper is the use of the velocity–pressure formulation: it keeps the divergence-free structure visible, reveals the flat ridge-ray geometry at least on short time intervals, and leads to an exact apex-dynamics reduction on the distinguished ridge rays. This reduction produces an explicit finite-time blow-up mechanism at the ridge apex $(x,\xi)=(0,\pm 1)$.
The paper has three main outputs. First, we identify the special ridge rays on which (E1) reduces to a convection-free (1+1)(1+1)D reaction system of Constantin–Lax–Majda type, and we record the resulting exact apex blow-up mechanism for $t\in[0,T]$. Second, we derive the exact remainder equations around a prescribed background and prove singular weighted linear estimates for the remainder system. Third, we formulate a conditional nonlinear control principle in the spirit of Elgindi–Jeong: if a compatible background exists on $[0,T)$ with the coefficient bounds required by the weighted energy method, and if the remainder remains subordinate to the background singularity in the detecting norm, then the full solution inherits the same finite-time blow-up.
Our approach is complementary to the vorticity–stream and boundary-driven singularity frameworks in the recent literature. Here the analysis is carried out on the full reduced-plane geometry attached to the pressure–velocity reduction, with smooth functions and with symmetry replacing boundary/irregularity mechanisms in the handling of convection. What is unconditional in the present paper is the exact reduction from Boussinesq, the exact apex blow-up dynamics on the ridge, the apex flatness criterion at $x=0$, and the weighted remainder framework. What remains conditional is the construction of a full background away from the apex together with the rigidity properties needed to close the bootstrap without loss. In this sense, the manuscript isolates the background-extension problem as the main remaining step toward a complete nonlinear stability theorem for the blow-up scenario.
@@ 1. Introduction
The question of whether smooth solutions can develop a singularity in finite time is a central theme in nonlinear PDE and fluid mechanics. In this paper we study a $(1+2)$-dimensional closed subsystem (E1), rigorously derived from the inviscid 2D Boussinesq equations in velocity–pressure form under a parity ansatz. Our aim is twofold: first, to identify an explicit finite-time blow-up mechanism at the ridge apex; second, to formulate a perturbative stability theory around compatible backgrounds that is mathematically solid at the linear level and explicit about the remaining nonlinear obstruction.
A central advantage of the pressure–velocity formulation is that the divergence-free condition remains visible throughout the reduction, the flat ridge-ray structure is revealed directly at least for short times, and the apex dynamics on the distinguished rays can be isolated exactly on the full reduced-plane geometry. This point of view is complementary to formulations based on vorticity–stream variables. In the present paper, convection control is handled through the inherited symmetry of the reduced equations rather than through a physical boundary or a lower-regularity functional class. The resulting framework is therefore best described as a complementary pressure–velocity and smooth-function approach to finite-time blow-up for structures derived from the Boussinesq system.
@@
As emphasized by Elgindi–Jeong [17] and Elgindi–Pasqualotto [19], singularity formation for the inviscid 2D Boussinesq and 3D Euler systems has a long history. We recall only a small selection of representative references here, emphasizing works most closely aligned with the blow-up mechanism and stability framework developed below. For the inviscid 2D Boussinesq system, classical local theory and conditional breakdown criteria go back to Cannon–DiBenedetto [12], Chae–Nam [3] (see also Chae–Kim–Nam [7]), and Taniuchi [29]; see also Wu’s lecture notes [30] and the Euler perspective of Beale–Kato–Majda [1] and Constantin [10, 11]. On the modeling side, explicit/didactic mechanisms include Chae–Constantin–Wu [2] and the CLM/De Gregorio lineage [9, 16], as well as 1D reductions for Boussinesq-type dynamics such as Choi–Kiselev–Yao [8]. For rigorous singularity constructions and perturbative stability scenarios in PDE settings, see Elgindi–Jeong [18, 17], Chen–Hou [4, 5], and the synthesis of Drivas–Elgindi [15]. In comparison with those works, the present paper adopts a different viewpoint: it starts from the pressure–velocity form, works with smooth functions on the full reduced-plane geometry associated with the symmetry reduction, and exploits symmetry rather than boundary effects or lower-regularity singular norms in the handling of convection. The benefit of this viewpoint is that the divergence-free structure and the ridge-ray geometry remain directly visible in the reduced equations, which in turn makes the exact apex-dynamics reduction transparent.
@@
    We identify the ridge rays on which (E1) reduces to a convection-free (1+1)(1+1)D reaction system of Constantin–Lax–Majda type, and we record the resulting exact finite-time blow-up profile at the ridge apex $(x,\xi)=(0,\pm 1)$ for $t\in[0,T]$.
@@
    We show that the pressure–velocity form reveals the divergence-free structure and the short-time flat ridge-ray geometry in a way that is compatible with the exact apex reduction.
@@
    We derive the exact remainder equations around a prescribed background in the $(x,\xi)$ variables, with all pure-background contributions retained in the background system.
@@
    We prove weighted singular linear estimates and formulate a conditional nonlinear remainder theorem of Elgindi type: once a compatible background is available with the required coefficient bounds, blow-up transfers from its apex dynamics to the full solution.
  • (5)
    We isolate the remaining open step in the program, namely the construction and control of a full background away from the apex together with the compatibility structure needed to close the nonlinear bootstrap.
Organization. Section 2 derives the closed subsystem (2.6) from the inviscid 2D Boussinesq equations and identifies the ridge rays on which (E1) reduces to a (1+1)(1+1)D CLM-type reaction system. Section 2.6 rewrites the background and remainder system in the $(x,\xi)$ variables. Section 4.4 records the exact remainder equations, and Section 4 collects the candidate background formulas, the explicit apex dynamics, and the corresponding weighted pointwise and energy bounds used in the conditional framework. Section 5 proves the singular weighted linear estimates, formulates the conditional nonlinear remainder mechanism, and explains how blow-up transfers from the background to the full solution once the remaining compatibility and background hypotheses are verified.
@@ 2.4. Background
$\left\{\begin{aligned} \lambda U_{t}-\tfrac{1}{2}V\,U&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xU_{x}-(G+V)\xi U_{\xi},\\[5.69054pt] \lambda V_{t}-V^{2}+U^{2}&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xV_{x}-(G+V)\xi V_{\xi}+\tfrac{1}{x}P_{x}-\tfrac{1+\xi^{2}}{x^{2}}\xi{P_{\xi}},\\[5.69054pt] \lambda G_{t}+G^{2}&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xG_{x}-\tfrac{1}{2}\xi(G^{2}+V^{2})_{\xi}-\tfrac{\mu}{x}P_{x}-\tfrac{\mu(1+\xi^{2})}{x^{2}\xi^{2}}\xi{P_{\xi}},\\[5.69054pt] G-V&=-\tfrac{1}{1+\xi^{2}}(\xi^{2}xG_{x}-xV_{x})-\xi(G_{\xi}+V_{\xi}).\end{aligned}\right$. (2.11)
The Poisson pressure equation (2.7) becomes
$\left\{\begin{aligned} \Delta P:&=P_{xx}+\tfrac{1}{x}P_{x}+\tfrac{2}{x^{2}}\left(\xi^{2}+1\right)\xi P_{\xi}+\tfrac{1}{x^{2}}\left(\xi^{2}+1\right)^{2}P_{\xi\xi}\\ &=U^{2}-\tfrac{2}{x}(\xi^{2}+1)^{1/2}\xi(U^{2})_{\xi}+(\xi^{2}+1)^{-1/2}(U^{2})_{x}.\end{aligned}\right$. (2.12)
@@ 2.5. Ridge ray and Ridge functions
$\lambda=\tfrac{3}{2},\quad\mu=\tfrac{5}{3},$ (2.15)
@@
and a simplified Poisson equation for $P(t,x,\xi_{0})$.
$P_{x}$ $=\tfrac{3x}{8}\left(U^{2}-2V^{2}\right),$ (2.17)
@@
Separating $P_{x}$ from $V_{t}$ yields  (2.16) and eq:ridge-dynamics-2. This proves Claim (C) and completes the Proof of theorem 2.5. ∎
@@
Denote $\phi(\xi,\xi_{0})$ the famous end-vanishing smooth function on the inverval (refs needed here)
$\phi(\xi)=\exp\bigl(-(\xi^{2}-1)^{-2}\bigr),\quad\xi\in[-1,1].$ (2.21)
The function $\phi(\xi)$ has following nice properties:
[…]
Then we have the following:
Lemma 2.6.
@@
$\left\{\begin{aligned} &r=x^{2}+A_{1}\phi(\xi),\\ &a=a(x,\xi)=\frac{A}{\left(1+r\right)^{3}},\quad A>0,\\ &b=b(x,\xi)=\frac{Br}{\left(1+r\right)^{6}},\quad B>0.\end{aligned}\right$. (2.23)
@@
Further more
$\left\{\begin{aligned} &\partial_{\xi}^{k}r\bigl|_{\xi=\pm 1}=0,\quad k\in\mathbb{N}\\ &\partial_{\xi}^{k}a\bigl|_{\xi=\pm 1}=0,\\ &\partial_{\xi}^{k}b\bigl|_{\xi=\pm 1}=0,\\ &\partial_{\xi}^{k}U\bigl|_{\xi=\pm 1}=0,\\ &\partial_{\xi}^{k}V\bigl|_{\xi=\pm 1}=0,\\ &\partial_{\xi}^{k}G\bigl|_{\xi=\pm 1}=0.\end{aligned}\right$. (2.25)
Proof of 2.6.
Since each $\xi$-derivative will introduce a factor of $\phi$ or turn $\partial_{\xi}^{k}\phi$ into $\partial_{\xi}^{k+1}\phi$, the claims in (2.25) follow from (2.22). ∎
@@ 2.6. Ridge blow-up and the full-wedge extension problem
The preceding theorem gives an explicit blowing-up ridge core, and in particular an explicit apex blow-up mechanism at $(x,\xi)=(0,\pm 1)$. The full background problem is to extend that ridge core to a classical wedge solution of (2.11). The next theorem isolates the apex flatness mechanism at $x=0$, while the full extension away from the apex remains a separate issue.
Theorem 2.8 (Preservation of ridge flatness at $x=0$).
@@
Then the ridge flatness at $x=0$ is preserved by the PDE system (2.11).
$\left\{\begin{aligned} &\lim_{x\to 0}\bigl((\partial_{\xi}U)_{t}(t,x,\xi)\bigr|_{\xi=\pm 1}\bigr)=0,\\ &\lim_{x\to 0}\bigl((\partial_{\xi}V)_{t}(t,x,\xi)\bigr|_{\xi=\pm 1}\bigr)=0,\\ &\lim_{x\to 0}\bigl((\partial_{\xi}G)_{t}(t,x,\xi)\bigr|_{\xi=\pm 1}\bigr)=0,\end{aligned}\right$. (2.27)
and
$\left\{\begin{aligned} &\lim_{x\to 0}\bigl((\partial_{\xi}P)_{t}(t,x,\xi)\bigr|_{\xi=\pm 1}\bigr)=0.\end{aligned}\right$. (2.28)
Proof of Preservation of ridge flatness at $x=0$ 2.8.
Differentiation of the $U$-equation in (2.11) with respect to $\xi$ leads to.
@@
$\left\{\begin{aligned} \lambda(U_{\xi})_{t}&=-xU_{x}\tfrac{2\xi}{\left(\xi^{2}+1\right)^{2}}(G+V)\\ &+H_{1}(F,xF_{x})\cdot J_{1}(F_{\xi\xi},F_{x\xi},F_{\xi})\quad F\in\{U,V,G,P\}.\end{aligned}\right$. (2.29)
Differentiation of the $V$-equation in (2.11) with respect to $\xi$ leads to.
$\left\{\begin{aligned} \lambda(V_{\xi})_{t}&=-xV_{x}\tfrac{2\xi}{\left(\xi^{2}+1\right)^{2}}(G+V)\\ &+H_{2}(F,xF_{x})\cdot J_{2}(F_{\xi\xi},F_{x\xi},F_{\xi})\quad F\in\{U,V,G,P\}.\end{aligned}\right$. (2.30)
Differentiation of the $G$-equation in (2.11) with respect to $\xi$ leads to.
$\left\{\begin{aligned} \lambda(G_{\xi})_{t}&=-xG_{x}\tfrac{2\xi}{\left(\xi^{2}+1\right)^{2}}(G+V)\\ &+H_{3}(F,xF_{x})\cdot J_{3}(F_{\xi\xi},F_{x\xi},F_{\xi})\quad F\in\{U,V,G,P\}.\end{aligned}\right$. (2.31)
Here $H_{1},H_{2},H_{3}$ denote three vector functions whose elements are linear and homogeneous function of its arguments. Similarly $J_{1},J_{2},J_{3}$ denote three vector functions whose elements are linear and homogeneous function of its arguments. Since the arguments of the latter three are $\xi$-derivatives of extremely flat ridge functions, they vanish on the ridges (cf.(2.25)).
The claims in (2.27) follow after set $\xi=\pm 1$ in (2.29),(2.30),(2.31), and then take the limit of $x\to 0$.
From (2.12), we obtain
$\left\{\begin{aligned} &P_{\xi\xi}+\tfrac{2\xi}{\xi^{2}+1}P_{\xi}\\ &=-\tfrac{x\left(P_{x}+xP_{xx}\right)}{\left(\xi^{2}+1\right)^{2}}+\tfrac{2x\left(xU_{x}-\left(\xi^{3}+\xi\right)U_{\xi}\right)U(\tau,x,\xi)}{\left(\xi^{2}+1\right)^{5/2}}\\ &\quad+\tfrac{x^{2}U^{2}}{\left(\xi^{2}+1\right)^{2}}\end{aligned}\right$. (2.32)
[…]
$P_{t\xi\xi}+\tfrac{2\xi}{\xi^{2}+1}P_{t\xi}=0,$ (2.33)
Thus the solution is
$\partial_{t}(P_{\xi})(t,0,\xi)=\tfrac{C(t)}{\xi^{2}+1},$ (2.34)
Since $(P_{\xi})(0,0,\xi)=0$, we have to set $c(t)=0$ and obtain $\partial_{t}(P_{\xi})(t,0,\xi)=0$. This completes the proof of Theorem2.8.
Remark 2.9 (Interpretation of Theorem 2.8).
If Theorem 2.8 is taken as established, then the explicit ridge reduction continues to govern the apex dynamics at $(x,\xi)=(0,\pm 1)$: the first-order ridge-flatness constraints are propagated at the apex, so the pointwise ODE system (2.16) remains the correct leading-order mechanism for the blow-up there. What this theorem does not give by itself is a closed-form description of the full background for general $x>0$ and $|\xi|\leq 1$. Accordingly, the explicit part of the present theory is the apex blow-up mechanism, while the extension away from the apex remains conditional and is delegated to the background/control problem for the full wedge.
Conjecture 2.10 (Existence of a wedge background blow-up profile).
There exist a time $T>0$ and smooth functions
$U,V,G,P\in C^{\infty}\bigl([0,T)\times\Omega\bigr)$
with smooth initial data chosen compatibly with the ridge/apex symmetry class of the paper, such that the following hold.
(1) Wedge background evolution. The quadruple $(U,V,G,P)$ solves the background system (2.11) on $0,T)\times\Omega$ with the required compatibility, regularity, and first-order ridge-flatness conditions at the apex points $(x,\xi)=(0,\pm 1)$.
(2) Preservation of the apex blow-up dynamics. Along the apex trajectories $(x,\xi)=(0,\pm 1)$, the solution reduces to the distinguished ridge ODE dynamics identified in Theorem 2.7, and the corresponding finite-time blow-up law is preserved up to time $T$. In particular,
$|V(t,0,\pm 1)|\sim\frac{c_{*}}{T-t}\qquad\text{as }t\uparrow T,$
for some constant $c_{*}>0$.
(3) Critical-size background bounds. There exist constants $C_{1},C_{2}>0$ such that for all $t\in[0,T)$,
$\|V(t)\|_{L^{\infty}(\Omega)}\leq\frac{C_{1}}{T-t},\qquad\|U(t)\|_{L^{\infty}(\Omega)}\leq\frac{C_{2}}{T-t}.$
Moreover, the associated fields $G$ and the derivatives of $(U,V,G,P)$ required by the perturbative analysis obey bounds of the same critical scale.
(4) Compatibility with the perturbative bootstrap. The background coefficient bounds are strong enough to justify all weighted estimates used in the perturbation theory, so that the bootstrap assumptions can in principle be closed up to time $T$ for sufficiently small perturbations of this background.